# Forecasting CPI inflation under economic policy and geo-political uncertainties

### v1

@@ 2 Data and Preliminary Analysis
This work is premised on forecasting the monthly CPI inflation series for the four countries: Brazil, Russia, India, and China based on their historical data and various uncertainty indices EPU and GPRC [^2] Two news-based measures of uncertainty are used in this study: the geopolitical risk uncertainty index (GPRC) [24] and the economic policy uncertainty index (EPU) of [19]. EPU stands for Economic Policy Uncertainty, which represents the likelihood of changes in monetary, fiscal, or regulatory policies that can impact the decision-making behavior of investors and consumers [19]. The geopolitical risk index[24] measures the occurrence of geopolitical events such as terrorism, local and regional political instability, political violence, coups d’état, territorial conflicts, and war. These events contribute to the uncertainty in the geopolitical landscape. Please refer to http://www.policyuncertainty.com/ and https://www.matteoiacoviello.com/gpr.htm for more information on the EPU and GPRC index calculations.. To evaluate the performance of our proposed method, we consider two forecast horizons of length 12 months and 24 months for each of the countries under study. In this section, we discuss the macroeconomic data considered in our analysis and their characteristics, along with the data preparation step (Section 2.1) and the causality analyses (Section 2.2) performed in this study. Furthermore, the implementable codes and the datasets used in this study are made available on GitHub repository for the reproducibility of the results.
@@ 2.1 Data Preparation and Global Charateristics
[…]
$\pi_{\text{Jan }}^{2003}=\left[\frac{CPI_{Jan}^{2003}-CPI_{\text{Jan }}^{2002}}{CPI_{Jan}^{2002}}\right]*100\%,$
@@
where $\pi_{Jan}^{2003}$ denotes the derived CPI inflation number for 2003-Jan, calculated using CPI values of 2003-Jan and 2002-Jan (prior year and same month). In addition to the CPI inflation, we also consider the impact of two uncertainty measures, namely the EPU and GPRC index, on the future projections of CPI inflation numbers for the BRIC countries. As part of the data pre-processing step, we perform the log transformation on the EPU series to stabilize its variance. Sequentially, we decompose the CPI inflation, log-transformed EPU, and GPRC series using the Hodrick and Prescott (HP) filter [49] and the Christiano-Fitzgerald (CF) random walk filter [50]. These economic filters play a vital role in extracting the trends and business cycles associated with the time series. Further details on the HP filter and CF filter are provided in Appendix 7.2.
@@
In the analysis, we study the global features of the CPI inflation series, log-transformed EPU series, and GPRC series to identify their structural patterns and detect appropriate techniques for long-term forecasting. Primarily we focused on six popular time series characteristics, including skewness, kurtosis, non-linearity, long-range dependence, seasonality, and stationarity [51]. To assess the long-range dependency of the time series, we utilize the Hurst exponent. To analyze the non-linear characteristics of the series, we employ Tsay’s test and Keenan’s one-degree test for non-linearity. To determine the stationarity of the series, we conduct the Kwiatkowski–Phillips–Schmidt–Shin test. Lastly, to identify any seasonal patterns in the series, we apply the Ollech and Webel’s test. The results of the statistical features, summarized in Table 1, depict that most of the macroeconomic time series are non-stationary except for the GPRC series of Brazil and the CPI inflation data of China. The majority of these series are positively skewed except for the log-transformed EPU series of Russia. Furthermore, most of the economic time series exhibit non-linear patterns which plays a crucial role in selecting the appropriate modeling approach for the dataset. Additionally, based on Ollech and Webel’s seasonality test and the Hurst exponent we observe that all the series are non-seasonal and are characterized by long-range dependency. Moreover, the ACF and PACF plots of the CPI inflation series (presented in Table 2) depict the presence of serial autocorrelation. These inherent patterns of the CPI inflation series motivate the design and application of hybrid forecasting approaches for long-term forecasting.
@@
Countries Series Skewness Kurtosis Non-Linearity Long-Range Dependence Seasonality Stationarity
Brazil CPI Inflation 1.67 3.57 Non-linear 0.73 Non-seasonal Non-stationary
log(EPU) -0.9 0.29 Non-linear 0.77 Non-seasonal Non-stationary
GPRC 2.06 6.60 Non-linear 0.63 Non-seasonal Stationary
Russia CPI Inflation 0.28 -0.98 Linear 0.78 Non-seasonal Non-stationary
log(EPU) -0.03 -0.12 Non-linear 0.79 Non-seasonal Non-stationary
GPRC 1.41 2.46 Linear 0.79 Non-seasonal Non-stationary
India CPI Inflation 0.71 0.17 Linear 0.82 Non-seasonal Non-stationary
log(EPU) 0.12 -0.47 Linear 0.80 Non-seasonal Non-stationary
GPRC 2.51 15.22 Non-linear 0.73 Non-seasonal Non-stationary
China CPI Inflation 0.82 1.08 Non-linear 0.69 Non Seasonal Stationary
log(EPU) 0.23 -0.87 Non-linear 0.82 Non-seasonal Non-stationary
GPRC 1.28 1.51 Non-linear 0.80 Non-seasonal Non-stationary
Table 2: Training data for FEWNet including target series CPI inflation (blue) with corresponding ACF, PACF plots and exogenous variables log-transformed EPU (orange) and GPRC (grey) for the BRIC countries.
Country
Training data
ACF plot
PACF plot
Brazil [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Russia [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
India [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
China [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 2.2 Causality Analysis
This section explores the causal impact of log-transformed EPU and GPRC uncertainties on the CPI inflation series for the BRIC countries. Several statistical methods, such as the Granger causality test, cross-convergent mappings, and transfer entropy, have been suggested in the literature to examine the causal relationship between two variables [52]. Nevertheless, the tests rely on certain data-level assumptions that are not met in the macroeconomic series analyzed in this work. In order to address this constraint, we employ wavelet coherence analysis (WCA) to examine the causal relationship between the series [53]. The WCA methodology offers a valuable method for examining the interconnection and simultaneous movement of two non-stationary signals in the time-frequency domain.
Figure 1: Wavelet coherence analysis plots - CPI inflation, log-transformed EPU (top) and CPI inflation, GPRC (bottom) for BRIC Countries
In the WCA plots, demonstrated in Figure 1, the first row depicts the WCA results between CPI inflation and log-transformed EPU series, and the second row shows the same for CPI inflation and GPRC series. In each of the plots horizontal axis measures the time, and the vertical axis shows the frequency dimension. Wavelet coherence represents the regions in the time-frequency domain where the two signals co-vary. In the plot, the transition from warmer colors (red) to colder colors (blue) indicates regions from higher coherence to lower coherence between the series. The arrow in the WCA plots indicates the lag phase relations between the analyzed signals. A zero-phase difference would hint toward the co-movement of the signals on any particular scale. The direction of an arrow has a specific connotation, with the arrows pointing to the right (left) indicating that the signals are in phase (anti-phase), i.e., the two signals are moving in the same direction or vice versa. As evident from the plots, there exists a significant causal relation between CPI inflation, log-transformed EPU, and GPRC series. This in turn, justifies the choice of EPU and GPRC as exogenous factors in forecasting the CPI inflation series of BRIC countries.
@@ 3 Proposed FEWNet Model
The real-world Consumer Price Index (CPI) inflation series exhibit notable features, including non-stationarity, nonlinearity, and long-term dependence, as evidenced in Table 1. In order to accurately represent this non-uniform time series, we utilize MODWT decomposition using Daubechies orthogonal wavelets of length 2, namely the Haar filter. This transformation has the ability to identify the specific time and frequency components of signals, successfully distinguishing between long-range connections, non-stationarity, and patterns within the signals [54]. By utilizing the MODWT-based multi-resolution analysis (MRA) decomposition on the series $Y_{t}$, we derive a collection of $\mathcal{K}+1$ distinct random variables called wavelet and scaling coefficients, which are not correlated with each other. In our suggested methodology, we employ separate neural networks within an ensemble framework to represent $\mathcal{K}+1$ random variables, instead of employing the original CPI inflation series $Y_{t}$. These neural networks are specifically engineered and trained to effectively manage the intricacies linked to the wavelet coefficients. By employing this approach, we may efficiently tackle the non-linear nature observed in the Consumer Price Index (CPI) inflation data of the BRIC nations. In addition, in order to capture the temporal causal association between the input series and the exogenous variables ($X_{j,t},\;j=1,2,\ldots,6$), we include them in each of the $\mathcal{K}+1$ wavelet transformed series of $Y_{t}$ in their corresponding neural networks. The forecasts produced by these $\mathcal{K}+1$ neural networks are combined to determine the future dynamics of $Y_{t}$. The mathematical expression for the $m$-step ahead forecasts of $Y_{t}$, designated as $\hat{Y}_{\mathcal{N}+m}$, based on $\mathcal{N}$ past data, is as follows:
@@
$\begin{gathered}\hat{\delta}_{\tilde{k},\mathcal{N}+1}=\alpha_{0,\tilde{k}}+\sum_{i=1}^{q}\beta_{i,\tilde{k}}\sigma(\alpha_{i,\tilde{k}}+\beta^{{}^{\prime}}_{i,\tilde{k}}\vec{\delta}_{\tilde{k}}+\gamma^{{}^{\prime}}_{i,\tilde{k}}\underline{X});\;\tilde{k}=1,2,\ldots,\mathcal{K}\\ \hat{S}_{\mathcal{K},\mathcal{N}+1}=\alpha_{0}^{S}+\sum_{i=1}^{q}\beta_{i}^{S}\sigma(\alpha_{i}^{S}+{\beta^{S}_{i}}^{{}^{\prime}}\vec{S}_{\mathcal{K}}+{\gamma^{S}_{i}}^{{}^{\prime}}\underline{X}),\end{gathered}$
where $\vec{\delta}_{\tilde{k}}$ and $\vec{S}_{\mathcal{K}}$ denotes the $p-6$ lagged observation of the corresponding wavelet coefficients and $\underline{X}$ indicates a single lagged vector of exogenous variables, $\alpha_{0,\tilde{k}},\alpha_{0}^{S},\alpha_{i,\tilde{k}},\alpha_{i}^{S}$ are the bias terms, $\beta_{i,\tilde{k}},\beta_{i}^{S}$ are the connection weights between hidden and output layer, $\gamma^{{}^{\prime}}_{i,\tilde{k}},{\gamma^{S}_{i}}^{{}^{\prime}}$ are the weight vectors between the exogenous variables and the hidden layer, $\beta^{{}^{\prime}}_{i,\tilde{k}},{\beta^{S}_{i}}^{{}^{\prime}}$ are the connection weights between lagged wavelet series and the hidden layer, and $\sigma$ indicates the nonlinear activation function. Within our framework, we commence by assigning random initial values to the connection weights. Subsequently, we employ the gradient descent back-propagation approach to train these weights, as described in the work by [55]. The aforementioned technique produces a forecast that is one step ahead for each of the wavelets and scaling coefficients. In order to make forecasts for several future time steps, we incorporate the most recent predictions into the input layer instead of using the last observed value, and we continue this procedure recursively. Ultimately, we combine the predictions produced by each of the base models in an ensemble setup and derive the forecast for a for a desired time horizon.
@@
where $\mathcal{V}$ indicates the validation set and $Y_{t}$ and $\hat{Y}_{t}$ are the original series and forecast at time $t$. Moreover, to stabilize the learning rate of the neural network, restrict overfitting, and to reduce the run-time of the proposed FEWNet framework, we set the number of hidden nodes arranged in a single hidden layer as $q=\left[\frac{p+1}{2}\right]$ following [51, 57]. The suggested FEWNet model is illustrated in Figure 3 using a schematic approach. The target time series is initially subjected to a MODWT-based MRA, resulting in multiple degrees of transformation, as shown in the figure. The economic uncertainties and geopolitical risk index series are filtered along with the CPI series using the CF filter and HP filter to generate the corresponding cycle (green circles) and trend (yellow circles) respectively. These six exogenous variables are then provided with the lagged wavelet series in the input layer of each of the $\mathcal{K}+1$ neural networks. The connection weights establish the connections between each input set and the hidden nodes, represented as lines with arrows. The results from the hidden layer are processed in the output layer, ultimately producing the forecast for the next step from the interconnected network. Ultimately, the future dynamics of the CPI inflation series are formed by producing multi-step ahead forecasts in a recursive fashion and then combining them in an ensemble. The utilization of wavelet decomposition to reveal the complex variability of the transformed series, and the subsequent modeling of these fluctuations with diverse economic filtered uncertainties using an auto-regressive neural network with auxiliary information within an ensemble framework, provides a valid rationale for naming the proposed model as FEWNet. The implementation methodology of the proposed FEWNet is outlined in Algorithm 1.
@@
Input: A target time series $\{Y_{t},\;t=1,2,\ldots\mathcal{N}\}$ indicating CPI inflation and exogenous series $X_{t}$ comprising of log-transformed EPU, and GPRC.
Output: $m\;(\geq 1)$ step-ahead forecast of the CPI inflation series $Y_{t}$.
@@
2 Determine the value of $\mathcal{K}$ as $\operatorname{log}_{e}(\mathcal{N})$, where $\mathcal{N}$ represents the size of the training dataset.

### v2

@@ 2 Data and Preliminary Analysis
This work is premised on forecasting the monthly CPI inflation series for the four countries: Brazil, Russia, India, and China, based on their historical data and various economic uncertainty measures such as EPU and GPRC. To evaluate the performance of our proposed method, we consider two forecast horizons, namely, a semi-long-term length of 12 months and a long-term length of 24 months, for each of the countries. This section details the datasets used in our analysis, their properties, the steps involved in data preparation (Section 2.1), and the causality analyses conducted (Section 2.2).
@@ 2.1 Data Preparation and Global Charateristics
In this study, we use monthly data on CPI inflation, EPU, and GPRC for BRIC countries from January 2003 (2003-01) to November 2021 (2021-11). For the semi-long-term forecasting task, we set the training period from 2003-01 to 2020-11 and the test horizon as 2020-12 to 2021-11, whereas for the long-term forecasting, our training data spans from 2003-01 to 2019-11 and test period extends from 2019-12 to 2021-11. Our target CPI inflation (also headline inflation) is measured as the year-on-year percentage change in the consumer price index (CPI) series obtained from the Federal Reserve Economic Data (FRED) https://fred.stlouisfed.org/. These numbers are seasonally unadjusted and are calculated based on the following approach:
$\pi_{\text{Jan }}^{2003}=\left[\frac{CPI_{Jan}^{2003}-CPI_{\text{Jan }}^{2002}}{CPI_{Jan}^{2002}}\right]*100\%,$ (1)
where $\pi_{Jan}^{2003}$ denotes the derived CPI inflation number for 2003-Jan, calculated using CPI values of 2003-Jan and 2002-Jan (prior year and same month).
We use two of the most popular news-based measures of uncertainty: the EPU index developed by Baker et al. (2016) and the GPRC index developed by Caldara & Iacoviello (2022). EPU represents the likelihood of changes in monetary, fiscal, or regulatory policies that can impact the decision-making behavior of investors and consumers Baker et al. (2016). According to Baker et al. (2016), the EPU Index is constructed by analyzing the frequency of certain terms in newspaper articles from a variety of sources, including the “Access World News Bank Service”. The index captures articles that mention key terms related to economic and policy uncertainty. Specifically, it searches for combinations of terms such as ‘economics’, ‘deficit’, ‘uncertainty’, ‘monetary policy’, ‘regulation’, ‘legislation’, ‘White House’, ‘Federal Reserve’, or ‘Congress’. This electronic text search process helps quantify the level of policy uncertainty by tracking the presence of these terms over time in major newspapers. The data on EPU is sourced from http://www.policyuncertainty.com/. Similarly, GPRC measures the occurrence of geopolitical events such as terrorism, local and regional political instability, political violence, coups, territorial conflicts, and war. These events contribute to the uncertainty in the geopolitical landscape. According to Caldara & Iacoviello (2022), the GPRC index is constructed by conducting electronic text searches of specific terms related to geopolitical tensions across 10 prominent daily newspapers. The GPRC index categorizes search terms into eight distinct categories to measure various facets of war-related events and geopolitical tensions: War Threats (1), Peace Threats (2), Military Buildups (3), Nuclear Threats (4), Terror Threats (5), Beginning of War (6), Escalation of War (7), Terror Acts (8). The data on GPRC is obtained from https://www.matteoiacoviello.com/gpr.htm.
@@
As part of the data pre-processing step, we perform the log transformation on the EPU series to stabilize its variance. Sequentially, we decompose the CPI inflation, log-transformed EPU, and GPRC series using the Hodrick and Prescott (HP) filter Hodrick & Prescott (1997) and the Christiano-Fitzgerald (CF) random walk filter Christiano & Fitzgerald (2003). These economic filters play a vital role in extracting the trends and business cycles associated with the time series. Further details on the HP filter and CF filter are provided in Appendix 8.3. The summary statistics of our variables under study are provided in Table 1.
Table 1: Summary statistics of the datasets used in this study
Countries Series Obs. Min. Value Max. Value ${Q}_{1}$ Median Mean ${Q}_{3}$ CoV Entropy
Brazil CPI Inflation 227 1.88 17.24 4.31 5.7 6.11 6.86 46.86 5.41
log(EPU) 227 1.35 2.83 2.00 2.15 2.16 2.3 10.94 5.65
GPRC 227 0.01 0.23 0.03 0.04 0.05 0.06 66.08 15.13
Russia CPI Inflation 227 2.2 16.93 5.46 7.61 8.44 11.60 46.01 5.38
log(EPU) 227 1.38 2.90 1.96 2.15 2.15 2.35 13.28 5.65
GPRC 227 0.22 2.23 0.42 0.58 0.67 0.83 51.46 6.05
India CPI Inflation 227 1.08 16.22 4.51 6.09 6.69 8.70 43.43 5.41
log(EPU) 227 1.37 2.45 1.72 1.90 1.90 2.06 11.70 5.68
GPRC 227 0.06 0.89 0.14 0.18 0.20 0.24 44.93 7.86
China CPI Inflation 227 -1.79 8.81 1.51 2.11 2.52 3.21 75.55 5.47
log(EPU) 227 1.42 2.99 1.96 2.18 2.24 2.51 16.97 5.63
GPRC 227 0.21 1.52 0.36 0.47 0.53 0.65 46.05 6.27
@@
In the analysis, we study the global features of the CPI inflation series, log-transformed EPU series, and GPRC series to identify their structural patterns and detect appropriate techniques for long-term forecasting. Primarily, we focused on seven popular time series characteristics, including skewness, kurtosis, non-linearity, long-range dependence, seasonality, stationarity, and outlier check Hyndman & Athanasopoulos (2018). We utilize the Hurst exponent to assess the long-range dependency of the time series. To analyze the non-linear characteristics of the series, we employ Tsay’s test and Keenan’s one-degree test for non-linearity. To determine the stationarity of the series, we conduct the Kwiatkowski–Phillips–Schmidt–Shin test. Lastly, we apply Ollech and Webel’s test to identify any seasonal patterns in the series. The results of the statistical features, summarized in Table 2, depict that most of the macroeconomic time series are non-stationary except for the GPRC series of Brazil and the CPI inflation data of China. The majority of these series are positively skewed except for the log-transformed EPU series of Russia. Furthermore, most of the economic time series exhibit non-linear patterns, which plays a crucial role in selecting the appropriate modeling approach for the dataset. Additionally, based on Ollech and Webel’s seasonality test and the Hurst exponent we observe that all the series are non-seasonal and are characterized by long-range dependency. Moreover, the ACF and PACF plots of the CPI inflation series (presented in Table 3) depict the presence of serial autocorrelation. These inherent patterns of the CPI inflation series motivate the design and application of hybrid forecasting approaches for long-term forecasting. We also performed the Bonferroni outlier test Weisberg & Cook (1982) based on studentized residuals. Our findings indicate that none of the CPI inflation series has significant outliers based on the adjusted p-values. However, while inflation remains relatively steady, policy uncertainty and geopolitical risks experience sudden and substantial fluctuations. This differential behavior underscores the necessity for policymakers to vigilantly monitor these indicators and implement strategic interventions aimed at mitigating the adverse effects of such outliers on the broader economy. We further examined the presence of structural breakpoints in the CPI inflation series to account for regime shifts and their potential impact on our downstream forecasting models, utilizing the ordinary least squares (OLS)-based CUSUM test Ploberger & Krämer (1992). By examining the cumulative sum of recursive residuals, the test can reveal deviations from the model’s stability, indicating potential breakpoints. In our study, the application of the OLS-based CUSUM test allows us to isolate and identify periods of structural change in the CPI inflation series. The CPI inflation series for these countries does not exhibit statistically significant breakpoints. The absence of statistically significant breakpoints in the CPI inflation series for Brazil, Russia, India, and China suggests that these series have remained relatively stable over the observed period. This stability implies that regime shifts are unlikely to pose significant challenges for forecasting models based on these series.
@@
Countries Series Skewness Kurtosis Non-Linearity Long-Range Seasonality Stationarity Outlier
Dependence
Brazil CPI Inflation 1.67 3.57 Non-linear 0.73 Non-seasonal Non-stationary No outliers detected
log(EPU) -0.9 0.29 Non-linear 0.77 Non-seasonal Non-stationary 1 outlier detected
GPRC 2.06 6.60 Non-linear 0.63 Non-seasonal Stationary 4 outliers detected
Russia CPI Inflation 0.28 -0.98 Linear 0.78 Non-seasonal Non-stationary No outliers detected
log(EPU) -0.03 -0.12 Non-linear 0.79 Non-seasonal Non-stationary No outliers detected
GPRC 1.41 2.46 Linear 0.79 Non-seasonal Non-stationary 2 outliers detected
India CPI Inflation 0.71 0.17 Linear 0.82 Non-seasonal Non-stationary No outliers detected
log(EPU) 0.12 -0.47 Linear 0.80 Non-seasonal Non-stationary No outliers detected
GPRC 2.51 15.22 Non-linear 0.73 Non-seasonal Non-stationary 1 outlier detected
China CPI Inflation 0.82 1.08 Non-linear 0.69 Non Seasonal Stationary No outliers detected
log(EPU) 0.23 -0.87 Non-linear 0.82 Non-seasonal Non-stationary No outliers detected
GPRC 1.28 1.51 Non-linear 0.80 Non-seasonal Non-stationary 1 outlier detected
Table 3: Training data for FEWNet including target series CPI inflation (blue), and exogenous variables log-transformed EPU (red) and GPRC (green). ACF, PACF, and OLS-based CUSUM test plots of CPI inflation series of the BRIC countries.
Training data (Jan 2003 to Nov 2019) ACF plot PACF plot OLS-based CUSUM test
Brazil
Russia
India
China
@@ 2.2 Causality Analysis
This section explores the causal impact of log-transformed EPU and GPRC uncertainties on the CPI inflation series for the BRIC countries. Several statistical methods, such as the Granger causality test, cross-convergent mappings, and transfer entropy, have been suggested in the literature to examine the causal relationship between two variables Pierce & Haugh (1977). Nevertheless, the tests rely on certain data-level assumptions that are not met in the macroeconomic series analyzed in this work. In order to address this constraint, we employ wavelet coherence analysis (WCA) to examine the causal relationship between the series Grinsted et al. (2004). The WCA methodology offers a valuable method for examining the interconnection and simultaneous movement of two non-stationary signals in the time-frequency domain.
Figure 1: Wavelet coherence analysis plots of CPI inflation with log-transformed EPU (left) and CPI inflation with GPRC (right) for (a) Brazil, (b) Russia, (c) India, and (d) China.
In the WCA plots, demonstrated in Figure 1, the first row depicts the WCA results between CPI inflation and log-transformed EPU series, and the second row shows the same for CPI inflation and GPRC series. In each of the plots, the horizontal axis measures the time, and the vertical axis shows the frequency dimension. Wavelet coherence represents the regions in the time-frequency domain where the two signals co-vary. In the plot, the transition from warmer colors (red) to colder colors (blue) indicates regions from higher coherence to lower coherence between the series. The arrow in the WCA plots indicates the lag phase relations between the analyzed signals. A zero-phase difference would hint toward the co-movement of the signals on any particular scale. The direction of an arrow has a specific connotation, with the arrows pointing to the right (left) indicating that the signals are in phase (anti-phase), i.e., the two signals are moving in the same direction or vice versa. As evident from the plots, there exists a significant causal relation between CPI inflation, log-transformed EPU, and GPRC series. This in turn, justifies the choice of EPU and GPRC as exogenous factors in forecasting the CPI inflation series of BRIC countries.
@@ 3 Proposed FEWNet Model
The real-world CPI inflation series exhibit notable features, including non-stationarity, non-linearity, and long-term dependence, as evidenced in Table 2. In order to accurately represent this non-uniform time series, we utilize MODWT decomposition using Daubechies orthogonal wavelets of length 2, namely the Haar filter. This transformation has the ability to identify the specific time and frequency components of signals, successfully distinguishing between long-range connections, non-stationarity, and patterns within the signals Percival & Mofjeld (1997). By utilizing the MODWT-based multi-resolution analysis (MRA) decomposition on the series $Y_{t}$, we derive a collection of $\mathcal{K}+1$ distinct random variables called wavelet and scaling coefficients, which are not correlated with each other. In our suggested methodology, we employ separate neural networks within an ensemble framework to represent $\mathcal{K}+1$ random variables instead of employing the original CPI inflation series $Y_{t}$. These neural networks are specifically engineered and trained to effectively manage the intricacies linked to the wavelet coefficients. By employing this approach, we may efficiently tackle the non-linear nature observed in the CPI inflation data of the BRIC nations. In addition, in order to capture the temporal causal association between the input series and the exogenous variables ($X_{j,t},\;j=1,2,\ldots,6$), we include them in each of the $\mathcal{K}+1$ wavelet transformed series of $Y_{t}$ in their corresponding neural networks. The forecasts produced by these $\mathcal{K}+1$ neural networks are combined to determine the future dynamics of $Y_{t}$. The mathematical expression for the $m$-step ahead forecasts of $Y_{t}$, designated as $\hat{Y}_{\mathcal{N}+m}$, based on $\mathcal{N}$ past data, is as follows:
@@
$\begin{gathered}\hat{\delta}_{\tilde{k},\mathcal{N}+1}=\alpha_{0,\tilde{k}}+\sum_{i=1}^{q}\beta_{i,\tilde{k}}\sigma(\alpha_{i,\tilde{k}}+\beta^{{}^{\prime}}_{i,\tilde{k}}\vec{\delta}_{\tilde{k}}+\gamma^{{}^{\prime}}_{i,\tilde{k}}\underline{X});\;\tilde{k}=1,2,\ldots,\mathcal{K}\\ \hat{S}_{\mathcal{K},\mathcal{N}+1}=\alpha_{0}^{S}+\sum_{i=1}^{q}\beta_{i}^{S}\sigma(\alpha_{i}^{S}+{\beta^{S}_{i}}^{{}^{\prime}}\vec{S}_{\mathcal{K}}+{\gamma^{S}_{i}}^{{}^{\prime}}\underline{X}),\end{gathered}$ (3)
where $\vec{\delta}_{\tilde{k}}$ and $\vec{S}_{\mathcal{K}}$ denotes the $p-6$ lagged observation of the corresponding wavelet coefficients and $\underline{X}$ indicates a single lagged vector of exogenous variables, $\alpha_{0,\tilde{k}},\alpha_{0}^{S},\alpha_{i,\tilde{k}},\alpha_{i}^{S}$ are the bias terms, $\beta_{i,\tilde{k}},\beta_{i}^{S}$ are the connection weights between hidden and output layer, $\gamma^{{}^{\prime}}_{i,\tilde{k}},{\gamma^{S}_{i}}^{{}^{\prime}}$ are the weight vectors between the exogenous variables and the hidden layer, $\beta^{{}^{\prime}}_{i,\tilde{k}},{\beta^{S}_{i}}^{{}^{\prime}}$ are the connection weights between lagged wavelet series and the hidden layer, and $\sigma$ indicates the non-linear activation function. Within our framework, we commence by assigning random initial values to the connection weights. Subsequently, we employ the gradient descent back-propagation approach to train these weights, as described in the work by Rumelhart et al. (1986). The aforementioned technique produces a forecast that is one step ahead for each of the wavelets and scaling coefficients. In order to make forecasts for several future time steps, we incorporate the most recent predictions into the input layer instead of using the last observed value, and we continue this procedure recursively. Ultimately, we combine the predictions produced by each of the base models in an ensemble setup and derive the forecast for a for a desired time horizon.
@@
where $\mathcal{V}$ indicates the validation set and $Y_{t}$ and $\hat{Y}_{t}$ are the original series and forecast at time $t$. Moreover, to stabilize the learning rate of the neural network, restrict overfitting, and to reduce the run-time of the proposed FEWNet framework, we set the number of hidden nodes arranged in a single hidden layer as $q=\left[\frac{p+1}{2}\right]$ following Hyndman & Athanasopoulos (2018); Panja et al. (2023). The suggested FEWNet model is illustrated in Figure 3 using a schematic approach. The target time series is initially subjected to a MODWT-based MRA, resulting in multiple degrees of transformation, as shown in the figure. The economic uncertainties and geopolitical risk index series are filtered along with the CPI inflation series using the CF filter and HP filter to generate the corresponding cycle (green circles) and trend (yellow circles), respectively. These six exogenous variables are then provided with the lagged wavelet series in the input layer of each of the $\mathcal{K}+1$ neural networks. The connection weights establish the connections between each input set and the hidden nodes, represented as lines with arrows. The results from the hidden layer are processed in the output layer, ultimately producing the forecast for the next step from the interconnected network. Ultimately, the future dynamics of the CPI inflation series are formed by producing multi-step ahead forecasts in a recursive fashion and then combining them in an ensemble. The utilization of wavelet decomposition to reveal the complex variability of the transformed series and the subsequent modeling of these fluctuations with diverse economic filtered uncertainties using an auto-regressive neural network with auxiliary information within an ensemble framework provides a valid rationale for naming the proposed model as FEWNet. The implementation methodology of the proposed FEWNet is outlined in Algorithm 1.
@@
Remark 1
Conventional machine learning and deep learning frameworks, such as XGBoost and TFT, often use a sliding window approach to transform time series forecasting into a supervised learning task; our proposed FEWNet architecture takes a distinct approach. We utilize an ensemble ARNNx framework on the MODWT decomposed training series and economically filtered exogenous variables. Unlike typical methods, the ARNNx model doesn’t reconstruct the economic series into an input-output supervised framework; instead, it leverages p-6 lagged observations (as we have 6 exogenous variables after HP and CF decomposition) of each of the decomposed training data and the previous lagged values of the exogenous variable to generate a one-step-ahead forecast using a non-linear function as given in Eq. (2) and (3). Furthermore, we iteratively update the training data with the latest forecast, developing a multi-step ahead forecast for each transformed series and aggregating forecasts from each wavelet decomposed series for our results. Our experimental setup involves two test horizons (24 months and 12 months) and distinct training periods corresponding to each of the horizons. For the 24 months period, we consider the training data from January 2003 – November 2019 and perform the decompositions of this available training dataset. For the 12-month horizon, the training period spans from January 2003 to November 2020, and the proposed FEWNet framework is fitted into this training set. Finally, for evaluation purposes, we use the original test data from each of these forecast horizons only to report the efficiency of the proposed FEWNet framework. Therefore, there is no data leakage while decomposing the data in this paper.
Input: The training data for the FEWNet model including the target time series $\{Y_{t},\;t=1,2,\ldots\mathcal{N}\}$ indicating CPI inflation and exogenous series $X_{t}$ comprising of log-transformed EPU, and GPRC.
Output: $\hat{Y}_{\mathcal{N}+1},\hat{Y}_{\mathcal{N}+2},\ldots,\hat{Y}_{\mathcal{N}+m};\;m\;(\geq 1)$ indicating $m$ step-ahead forecast of the CPI inflation series.
@@
2 Determine the value of $\mathcal{K}$ as $\lfloor\operatorname{log}_{e}(\mathcal{N})\rfloor$, where $\mathcal{N}$ represents the size of the training dataset.